• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name silwal

    Log entry time 16:04:57 on August30,2009

    Entry number 287103

    keyword=day shift summary

    8:00 LQ3 down, crygoneic trip
    8:50 No beam in all halls, RF problems
    9:35 Beam back on, back up to 60 uA
    9:50 Jack came in for LQ3 reset. But turned out that LQ3 was in a state to be reset without hall access. LQ3 reset, and is up now.
    10:11 Turned right arm det HV back on. Did not realize that it had been off all the while!
    10:47 requested a harp scan. beam at 5 uA
    10:35 MCC having issues with the harp software. The need some time. So beam back up to 60 uA.
    11:27 MCC ready for harp scan. beam at 5 uA.
    11:39 target moved to empty from hydrogen, raster off, MCC starting harp scan.
    11:52 harp done, moving back to hydrogen target.
    11:56 beam back up to 60 uA
    15:00 beam down on all halls. minor magnet part replacement. MCC gave us about 15 minutes before the beam is back up again.
    16:00 beam not back on yet.